# Cutoff for Contingency Table Random Walks

### v1

@@ 5.4 Cutoff for a Family of Random Walks on ${\mathbb{Z}}_{q}$
Let $\{X_{t}\}^{(n)}$ be an arbitrary family of irreducible symmetric continuous-time Markov chains on the torus ${\mathbb{Z}}_{q}^{n}$ with the setup in Sections 3.1 and 5.1. Specifically, we define the increment of each jump mapped to $({\mathbb{Z}}\cap[-q/2,q/2))^{n}$, $Y^{(n)}$, to have a correlation matrix $\Gamma_{n}$, and require $\|Y^{(n)}\|_{1}\leq r=r(n)$, and $\|Y^{(n)}\|_{\infty}<q/2$. Suppose that $Y^{(n)}$ has the same marginal variance by coordinate of $\dfrac{\sigma^{2}}{n}$, where $\sigma^{2}=\sigma^{2}(n)\leq r^{2}$ by the previous requirement. Let $\varepsilon\in(0,1)$ be arbitrary. Then we have the following,
@@
    Suppose that $\{Y\}^{(n)}$ meets the Decay Condition and $\{\Gamma_{n}\}$ meets the Correlation Condition with bounding function $g(\alpha)$. Let
@@
    On the other hand, as $n\to\infty$ and $q\to\infty$ with $\sup_{k\neq\ell}|\Gamma_{kl}|\ll\dfrac{1}{\log(n)}$, and $\dfrac{r^{2}}{q^{2}}\ll\dfrac{1}{\log(n)}$
@@
We see by Theorem 3 any family of Markov chains $\{X_{t}\}^{(n)}$ in the setup of Section 5.1 meeting the Decay and Correlation Conditions with $n\rightarrow\infty$, $q\rightarrow\infty$, $\dfrac{r^{2}}{q^{2}}\ll\dfrac{1}{\log(n)}$, and $\sup_{k\neq\ell}|\Gamma_{kl}|\ll\dfrac{1}{\log(n)}$ exhibits cutoff at the time $\dfrac{nq^{2}\log(n)}{4\pi^{2}\sigma^{2}}$, with a cutoff window of order $\dfrac{nq^{2}}{\sigma^{2}}$.
@@
Let $\{(S_{t}^{(n)})_{t\geq 0}\}_{n=1}^{\infty}$ be a family of symmetric and irreducible random walks on $\mathbb{Z}^{n}$ in continuous time so that each coordinate is equivariant with marginal variance $\dfrac{\sigma^{2}}{n}$, and $S_{t}^{(n)}$ meets the correlation condition. Then by the same reasoning as Example 3, we can set $q=q(n)$ large enough so that $S_{t}^{(n)}\mod q$ has cutoff at time $t=\dfrac{nq^{2}\log(n)}{4\pi^{2}\sigma^{2}}$.
6 $n\times n$ Contingency Tables
@@
We now consider the rate-$1$ continuous-time Diaconis-Gangolli random walk on $n\times n$ contingency tables over ${\mathbb{Z}}_{q}$, whose dynamics were introduced in Section 1.
@@
As in the $1\times n$ case, we will only keep track of the first $n-1$ rows and $n-1$ columns to gain irreducibility and aperiodicity. Using the analogous notations, we denote the resulting continuous-time process on ${\mathbb{Z}}_{q}^{(n-1)^{2}}$ by $\{X_{t}\}_{t\geq 0}$. Let $\{Y_{m}\}_{m\in{\mathbb{N}}}$ be the embedded discrete-time chain of $\{X_{t}\}_{t\geq 0}$, and assume without loss of generality that $X_{0}=Y_{0}=0$. Again, let $Y\in{\mathbb{Z}}^{(n-1)^{2}}$ be a lattice-valued random vector with the same law as that of $Y_{1}$ mapped to $({\mathbb{Z}}\cap[-q/2,q/2))^{(n-1)^{2}}$.
@@ 6.1 Inverting the Correlation Matrix
In this section, we present the inverse $\Gamma^{-1}\in\mathbb{R}^{(n-1)^{2}\times(n-1)^{2}}$ of the correlation matrix $\Gamma\in\mathbb{R}^{(n-1)^{2}\times(n-1)^{2}}$ of $Y$ for later use. Note that $\Gamma\in\mathbb{R}^{(n-1)^{2}\times(n-1)^{2}}$ is given by
$\Gamma_{(i,j),(i^{\prime},j^{\prime})}=\begin{cases}1&(i,j)=(i^{\prime},j^{\prime})\\ -\frac{1}{n-1}&i=i^{\prime}\text{ or }j=j^{\prime}\ \text{but not both}\\ \frac{1}{(n-1)^{2}}&i\neq i^{\prime}\text{ and }j\neq j^{\prime}\end{cases}.$
@@
We also note that the discrete-time increments have a covariance matrix of $\dfrac{4}{n^{2}}\Gamma$. We claim that the inverse $\Psi:=\Gamma^{-1}$ of $\Gamma$ is given by
$\Psi_{(i,j),(i^{\prime},j^{\prime})}=\begin{cases}\frac{4(n-1)^{2}}{n^{2}}&(i,j)=(i^{\prime},j^{\prime})\\ \frac{2(n-1)^{2}}{n^{2}}&i=i^{\prime}\text{ or }j=j^{\prime}\ \text{but not both}\\ \frac{(n-1)^{2}}{n^{2}}&i\neq i^{\prime}\text{ and }j\neq j^{\prime}\end{cases}.$
To verify $\Psi=\Gamma^{-1}$, we write
$\Gamma=\begin{bmatrix}\gamma_{(1,1)}&\gamma_{(1,2)}&\ldots&\gamma_{(n-1)^{2}}\end{bmatrix},\qquad\Psi=\begin{bmatrix}\psi_{(1,1)}&\psi_{(1,2)}&\ldots&\psi_{(n-1)^{2}}\end{bmatrix}$
and show $\langle\gamma_{(i,j)},\psi_{(i^{\prime},j^{\prime})}\rangle=\mathbbm{1}_{\{i=i^{\prime}\}}\mathbbm{1}_{\{j=j^{\prime}\}}$. For any fixed index $(i,j)$, there are exactly $2(n-2)$ indices sharing a row or column with the tuple $(i,j)$ and $(n-2)^{2}$ indices sharing neither a row or column with $(i,j)$. A direct computation then yields
$\langle\gamma_{(i,j)},\psi_{(i,j)}\rangle=\frac{(n-1)^{2}}{n^{2}}\left(4(1)+2(n-2)\frac{-2}{n-1}+\frac{(n-2)^{2}}{(n-1)^{2}}\right)=1.$
Next, suppose we have indices that share either a single row or a single column. Without loss of generality, we can consider $(i,j),\ (i^{\prime},j)$ with $i\neq i^{\prime}$, and the argument for $(i,j),\ (i,j^{\prime})$ with $j\neq j^{\prime}$ follows similarly. Let $u\odot v$ denote the element-wise multiplication between two vectors $u$ and $v$ of the same dimension, i.e. $u\odot v=(u_{1}v_{1},\dots,u_{n}v_{n})$ if $u,v$ are both of dimension $n$. By definition, we have
$\frac{n^{2}}{(n-1)^{2}}\left(\gamma_{(i,j)}\odot\psi_{(i^{\prime},j)}\right)_{(\ell,k)}=\begin{cases}-\frac{4}{n-1}&(\ell,k)=(i,j)\\ 2&(\ell,k)=(i^{\prime},j)\\ -\frac{2}{n-1}&\ell\notin\{i,i^{\prime}\}\text{ and }k=j\\ -\frac{1}{n-1}&\ell=i\text{ and }k\neq j\\ \frac{2}{(n-1)^{2}}&\ell=i^{\prime}\text{ and }k\neq j\\ \frac{1}{(n-1)^{2}}&\ell\notin\{i,i^{\prime}\}\text{ and }k\neq j\end{cases}.$
Since in the contingency table (with the last row and column ignored) there are $n-3$ rows not the $i$th or the $i^{\prime}$th and $n-2$ columns not the $j$th, it follows that
$\langle\gamma_{(i,j)},\psi_{(i^{\prime},j)}\rangle=\frac{(n-1)^{2}}{n^{2}}\left(\frac{-4}{n-1}+2-\frac{2(n-3)}{n-1}+(n-2)\left(\frac{2}{(n-1)^{2}}-\frac{1}{n-1}\right)+\frac{(n-3)(n-2)}{(n-1)^{2}}\right)=0.$
See Figure 1 below.
$(i^{\prime},j)$$(i,j)$$n-2$$n-3$
Figure 1: Diagram for indices with a shared column in the $n\times n$ contingency table with the last row and column ignored. Note that this diagram is only for counting purposes, and the rows and columns may have been permuted.
Now it remains to consider $\langle\gamma_{(i,j)},\psi_{(i^{\prime},j^{\prime})}\rangle$ such that $i\neq i^{\prime}$ and $j\neq j^{\prime}$. We can again write by definition that
$\frac{n^{2}}{(n-1)^{2}}\left(\gamma_{(i,j)}\odot\psi_{(i^{\prime},j^{\prime})}\right)_{(\ell,k)}=\begin{cases}1&(\ell,k)=(i,j)\\ \frac{4}{(n-1)^{2}}&(\ell,k)=(i^{\prime},j^{\prime})\\ -\frac{2}{n-1}&(\ell,k)\in\{(i,j^{\prime}),(i^{\prime},j)\}\\ -\frac{1}{n-1}&\left(\ell=i\text{ and }k\notin\{j,j^{\prime}\}\right)\text{ or}\left(\ell\notin\{i,i^{\prime}\}\text{ and }k=j\right)\\ \frac{2}{(n-1)^{2}}&\left(\ell=i^{\prime}\text{ and }k\notin\{j,j^{\prime}\}\right)\text{ or }\left(\ell\notin\{i,i^{\prime}\}\text{ and }k=j^{\prime}\right)\\ \frac{1}{(n-1)^{2}}&\ell\notin\{i,i^{\prime}\}\text{ and }k\notin\{j,j^{\prime}\}\end{cases}.$
Since in the contingency table (again with the last row and column ignored) there are $n-3$ rows not the $i$th or the $i^{\prime}$th and $n-3$ columns not the $j$th or the $j^{\prime}$th,
$\frac{n^{2}}{(n-1)^{2}}\langle\gamma_{(i,j)},\psi_{{(i^{\prime},j^{\prime})}}\rangle=1+\frac{4}{(n-1)^{2}}+2\left(-\frac{2}{n-1}\right)+2(n-3)\left(\frac{2}{(n-1)^{2}}-\frac{1}{n-1}\right)+\frac{(n-3)^{2}}{(n-1)^{2}}=0.$
See Figure 2 below. Hence, $\Gamma^{-1}=\Psi$ as claimed.
$(i^{\prime},j^{\prime})$$(i,j)$$n-3$$n-3$--
Figure 2: Diagram for indices that share no rows or columns in the $n\times n$ contingency table with the last row and column ignored. Note that this diagram is only for counting purposes, and the rows and columns may have been permuted.

### v2

@@ 5.4 Cutoff for a Family of Random Walks on ${\mathbb{Z}}_{q}^{n}$
Let $\{X_{t}\}^{(n)}$ be an arbitrary family of irreducible symmetric continuous-time Markov chains on the torus ${\mathbb{Z}}_{q}^{n}$ with the setup in Sections 3.1 and 5.1. Specifically, we define the increment of each jump mapped to $({\mathbb{Z}}\cap[-q/2,q/2))^{n}$, $Y^{(n)}$, to have a correlation matrix $\Gamma_{n}$, and require $\|Y^{(n)}\|_{1}\leq r=r(n)$, and $\|Y^{(n)}\|_{\infty}<q/2$. Suppose that $Y^{(n)}$ has the same marginal variance by coordinate of $\dfrac{\sigma^{2}}{n}$, where $\sigma^{2}=\sigma^{2}(n)$. Let $\varepsilon\in(0,1)$ be arbitrary. Then we have the following
@@
    Suppose that $\{Y^{(n)}\}$ meets the Decay Condition and $\{\Gamma_{n}\}$ meets the Correlation Condition with the bounding function $g(\alpha)$. Let
@@
    On the other hand, as $n\to\infty$ and $q\to\infty$ with $\sup_{k\neq\ell}|\Gamma_{n,k\ell}|\ll\dfrac{1}{\log(n)}$, and $\dfrac{r^{2}}{q^{2}}\ll\dfrac{1}{\log(n)}$
@@
We see by Theorem 3 any family of Markov chains $\{X_{t}\}^{(n)}$ in the setup of Section 5.1 meeting the Decay and Correlation Conditions with $n\rightarrow\infty$, $q\rightarrow\infty$, $\dfrac{r^{2}}{q^{2}}\ll\dfrac{1}{\log(n)}$, and $\sup_{k\neq\ell}|\Gamma_{n,k\ell}|\ll\dfrac{1}{\log(n)}$ exhibits cutoff at the time $\dfrac{nq^{2}\log(n)}{4\pi^{2}\sigma^{2}}$, with a cutoff window of order $\dfrac{nq^{2}}{\sigma^{2}}$.
@@
Let $\{(S_{t}^{(n)})_{t\geq 0}\}_{n=1}^{\infty}$ be a family of symmetric and irreducible random walks on $\mathbb{Z}^{n}$ in continuous time with bounded increments, such that each coordinate is equivariant with marginal variance $\dfrac{\sigma^{2}}{n}$, and $S_{t}^{(n)}$ meets the correlation condition. Then by the same reasoning as Example 3, we can set $q=q(n)$ large enough so that $S_{t}^{(n)}\mod q$ has cutoff at time $t=\dfrac{nq^{2}\log(n)}{4\pi^{2}\sigma^{2}}$, assuming $\sup_{k\neq\ell}|\Gamma_{n,k\ell}|\ll\dfrac{1}{\log(n)}$.
5.5 Discussion of the Regime $q\rightarrow\infty$
In the case of our main theorems, we require $q\gg n$ or $q\gg n^{2}$ for the upper bounds in order to show sufficient decay in the characteristic function of our lifted distribution $\Phi_{Y}$ to then bound the $\ell^{2}$ summation. This was done as a tool to gain control of the spectrum: they are sufficient (but not necessarily sharp) assumptions for the Decay Conditions to be met in the contingency table cases. In general, the requirement of the regime of $q$ depends on how one fulfills the Decay Condition for their specific model. For example, Nestoridi and Nguyen [25] showed cutoff for the Diaconis and Gangolli walk on $n\times n$ contingency tables over ${\mathbb{Z}}_{q}$ when $q$ is small compared to $n$, and [15] gives results for the simple random walk on ${\mathbb{Z}}_{q}^{n}$. Hermon and Olesker-Taylor in [15] also analyzed the cutoff times for random walks on the torus $\mathbb{Z}_{q}^{n}$ when $q$ is small.
For the mixing time lower bounds, the assumptions on $q$ are relatively mild and are primarily for the Taylor expansion analysis in the distinguishing statistics.
We do like to point out that there are cases where a fast-growing $q$ relative to $n$ is required for cutoff. In the regime where $q$ is small, we are not necessarily able to compare the rate of decay of the spectrum of our walk to the Gaussian characteristic function. Consider the following rate-$1$ random walk $X_{t}=(X_{t}^{(1)},X_{t}^{(2)},\dots,X_{t}^{(n)})$ on the torus ${\mathbb{Z}}_{q}^{n}$. For $j\geq 3$, each $X_{t}^{(j)}$ is a rate-$(1/n)$ simple random walk on ${\mathbb{Z}}_{q}$, independent of all the other coordinates, including $X_{t}^{(1)}$ and $X_{t}^{(2)}$. Each of $X_{t}^{(1)}$ and $X_{t}^{(2)}$ performs simple random walk jumps with rate $e^{-n}/n$, independent of each other and all the remaining n2n-2 coordinates. In addition, with rate $(1-e^{-n})/n$, $X_{t}^{(1)}$ and $X_{t}^{(2)}$ jump simultaneously but with independent simple random walk increments.
This dynamics has correlation matrix $I_{n}$, the $n\times n$ identity matrix, bounded increments, and is irreducible over the torus $\mathbb{Z}_{q}^{n}$. Since $X_{t}$ is symmetric, it converges to the uniform distribution $\pi$ on ${\mathbb{Z}}_{q}^{n}$. By Remark 5, as $n\to\infty$, there is a regime of $q=q(n)$ sufficiently large with respect to $n$, so that $X_{t}$ has cutoff at the time $\dfrac{nq^{2}\log(n)}{4\pi^{2}}$ with a cutoff window of order $\mathcal{O}(nq^{2})$.
The walk does not mix with cutoff in the regime when $n\rightarrow\infty$ and $q$ is even. In this regime, the difference of the first two coordinates $X_{t}^{(1)}$ and $X_{t}^{(2)}$ modulo $2$ is the primary bottleneck in the chain’s trajectory. To see this, let $D_{t}:=X_{t}^{(1)}-X_{t}^{(2)}\mod 2$, which is a rate-$(2e^{-n}/n)$ simple random walk on ${\mathbb{Z}}_{2}$ as a result of $q$ being even. We may assume $X_{0}=0$ due to transitivity of ${\mathbb{Z}}_{q}^{n}$, and let $H_{t}$ be the transition kernel of $X_{t}$. A direct calculation with the Kolmogorov forward and backward equations yields
$\|H_{t}(0,\cdot)-\pi\|_{\operatorname{TV}}\geq{\mathbb{P}}(D_{t}=0)-\frac{1}{2}=\frac{1}{2}\exp(-4te^{-n}/n).$
Thus, for $\varepsilon<1/2$,
$t^{(n)}_{\operatorname{mix}}(\varepsilon)\geq\frac{ne^{n}\log(\frac{1}{2\varepsilon})}{4}.$ (56)
We now give an upper bound to $t_{\operatorname{mix}}(1-\varepsilon)$ for $\varepsilon<1/q^{2}$. The following lemma can be proved directly using the optimal coupling definition of the total variation distance (see Proposition 4.7 of [18]).
Lemma 6.
Let $\mu_{1}$ and $\mu_{2}$ be probability measures supported on the discrete outcome set $\Omega$ with $\rVert\mu_{1}-\mu_{2}\rVert_{\operatorname{TV}}=\varepsilon_{1}$, and let $\nu_{1}$ and $\nu_{2}$ be probability measures supported on the discrete outcome set $\Omega^{\prime}$ with $\rVert\nu_{1}-\nu_{2}\rVert_{\operatorname{TV}}=\varepsilon_{2}$. The product measures $\mu_{1}\times\nu_{1}$ and $\mu_{2}\times\nu_{2}$ satisfy
$\rVert\mu_{1}\times\nu_{1}-\mu_{2}\times\nu_{2}\rVert_{\operatorname{TV}}\leq\varepsilon_{1}+\varepsilon_{2}-\varepsilon_{1}\varepsilon_{2}.$
We will view $X_{t}$ as the product chain of $(X_{1},X_{2})$ and each of the remaining coordinates. It is straightforward to show that at any $t\geq 0$, the law of $(X_{t}^{(1)},X_{t}^{(2)})$ has total variation distance at most $1-1/q^{2}$ with the uniform distribution of ${\mathbb{Z}}_{q}^{2}$. For each $j=3,\dots,n$, $X_{t}^{(j)}$ is a rate-$1/n$ simple random walk on ${\mathbb{Z}}_{q}$. It is well-known that the spectral gap of the discrete-time transition matrix of $X_{t}^{(j)}$ is $\gamma:=1-\cos(2\pi/q)$ (see, for example, 12.3.1 in [18]). Then, by Theorem 20.6 in [18], the total variance distance between the transition kernel of $X_{t}^{(j)}$ at time $t$ and the uniform distribution on ${\mathbb{Z}}_{q}$ is at most
$q\cdot\exp(-\gamma t/n).$
Therefore, at time
$t^{\star}=\frac{n\log(nq(1/q^{2}-\varepsilon)^{-1})}{1-\cos(2\pi/q)},$
the total variance distance between the transition kernel of $X_{t}^{(j)}$ and the uniform distribution on ${\mathbb{Z}}_{q}$ is at most $\frac{1/q^{2}-\varepsilon}{n}$. Now by Lemma 6,
$d(t^{\star})\leq 1-\frac{1}{q^{2}}+\sum_{j=3}^{n}\frac{1/q^{2}-\varepsilon}{n}\leq 1-\varepsilon.$
It follows that $t_{\operatorname{mix}}(1-\varepsilon)\leq t^{\star}$. Combining this with (56), we have that for $\varepsilon<1/q^{2}$
$\frac{t^{(n)}_{\operatorname{mix}}(\varepsilon)}{t^{(n)}_{\operatorname{mix}}(1-\varepsilon)}\geq\frac{e^{n}\log(\frac{1}{2\varepsilon})(1-\cos(\frac{2\pi}{q}))}{4\log(nq(1/q^{2}-\varepsilon)^{-1})}\to\infty$
as $n\to\infty$, which shows that $X_{t}$ does have cutoff with a fixed even $q$. Note that $X_{t}$ does not even exhibit pre-cutoff, which requires
$\sup_{0<\varepsilon<1/2}\limsup_{n\rightarrow\infty}\frac{t^{(n)}_{\operatorname{mix}}(\varepsilon)}{t^{(n)}_{\operatorname{mix}}(1-\varepsilon)}<\infty.$
6 $m\times n$ Contingency Tables
@@
We now consider the rate-$1$ continuous-time Diaconis-Gangolli random walk on $m\times n$ contingency tables over ${\mathbb{Z}}_{q}$, whose dynamics were introduced in Section 1. Again, we assume $\kappa_{1}n\leq m\leq\kappa_{2}n$ with constants $\kappa_{1},\kappa_{2}>0$. And WLOG we assume $\kappa_{1}n$ and $\kappa_{2}n$ are integers.
@@
As in the $1\times n$ case, we will only keep track of the first m1m-1 rows and n1n-1 columns to gain irreducibility and aperiodicity. Using the analogous notations, we denote the resulting continuous-time process on ${\mathbb{Z}}_{q}^{(m-1)(n-1)}$ by $\{X_{t}\}_{t\geq 0}$. Let $\{Y_{s}\}_{s\in{\mathbb{N}}}$ be the embedded discrete-time chain of $\{X_{t}\}_{t\geq 0}$, and assume without loss of generality that $X_{0}=Y_{0}=0$. Again, let $Y\in{\mathbb{Z}}^{(n-1)^{2}}$ be a lattice-valued random vector with the same law as that of $Y_{1}$ mapped to $({\mathbb{Z}}\cap[-q/2,q/2))^{(m-1)(n-1)}$.
Our proof strategy will be similar to sections 4 and 5. We will show that the Diaconis Gangolli walk on $m\times n$ contingency tables over ${\mathbb{Z}}_{q}$ meets the framework of our generalization in under the asymptotics where $n\rightarrow\infty$, $m\asymp n$, and $q\gg n^{2}$. In particular, we need to show the correlation and decay conditions hold. In Section 6.1, we diagonalize and invert the correlation matrix for the walk’s natural lift to $\mathbb{Z}^{(m-1)\times(n-1)}$. In section 6.2, we work directly with the characteristic function of the walk’s natural lift to show the decay condition. In Section 6.3, we verify that the correlation condition additionally holds by analyzing Schur complements of $\Gamma^{-1}$.
@@ 6.1 Inverting the Correlation Matrix
In this subsection, we present the inverse $\Gamma^{-1}\in\mathbb{R}^{(m-1)(n-1)\times(m-1)(n-1)}$ of the correlation matrix $\Gamma\in\mathbb{R}^{(m-1)(n-1)\times(m-1)(n-1)}$ of $Y$ for later use. Note that $\Gamma$ is given by
$\Gamma_{(i,j),(i^{\prime},j^{\prime})}=\begin{cases}1&(i,j)=(i^{\prime},j^{\prime})\\ -\frac{1}{n-1}&i=i^{\prime}\text{ and }j\neq j^{\prime}\\ -\frac{1}{m-1}&i\neq i^{\prime}\text{ and }j=j^{\prime}\\ \frac{1}{(m-1)(n-1)}&i\neq i^{\prime}\text{ and }j\neq j^{\prime}\end{cases}.$
@@
Also, $Y$ has covariance matrix $\dfrac{4}{mn}\Gamma$. We claim that the inverse $\Psi:=\Gamma^{-1}$ of $\Gamma$ is given by
$\Psi_{(i,j),(i^{\prime},j^{\prime})}=\begin{cases}\frac{4(m-1)(n-1)}{mn}&(i,j)=(i^{\prime},j^{\prime})\\ \frac{2(m-1)(n-1)}{mn}&i=i^{\prime}\text{ or }j=j^{\prime}\ \text{but not both}\\ \frac{(m-1)(n-1)}{mn}&i\neq i^{\prime}\text{ and }j\neq j^{\prime}\end{cases}.$
To verify the claim, we will give the eigenvalues and eigenvectors of $\Gamma$ and $\Psi$. First, the all-one vector $\mathbbm{1}_{(m-1)(n-1)}$ is an eigenvector of eigenvalue $\dfrac{1}{(m-1)(n-1)}$ for $\Gamma$ and of eigenvalue (m1)(n1)(m-1)(n-1) for $\Psi$. Next, if we define
$\delta_{i.}(\ell,k):=\begin{cases}-1&\ell=1\\ 1&\ell=i\\ 0&\ell\notin\{1,i\}\end{cases},\hskip 25.0pt\delta_{.j}(\ell,k):=\begin{cases}-1&k=1\\ 1&k=j\\ 0&k\notin\{1,j\}\end{cases},$
then for $2\leq i\leq m-1$, each $\delta_{i.}$ is an eigenvector for $\Gamma$ with eigenvalue $\dfrac{m}{(m-1)(n-1)}$, and eigenvector for $\Psi$ with eigenvalue $\dfrac{(m-1)(n-1)}{m}$. Similarly, one can show that for $2\leq j\leq n-1$, each $\delta_{.j}$ is an eigenvector for $\Gamma$ with eigenvalue $\dfrac{n}{(m-1)(n-1)}$ and an eigenvector for $\Psi$ with eigenvalue $\dfrac{(m-1)(n-1)}{n}$. Lastly, if we define the vectors
[…]
then each $\delta_{ij}$ is an eigenvector for eigenvalue $\dfrac{mn}{(m-1)(n-1)}$ for $\Gamma$, and an eigenvector for eigenvalue $\dfrac{(m-1)(n-1)}{mn}$ for $\Psi$, when $2\leq i\leq m-1$ and $2\leq j\leq n-1$. Hence, using the spectral decomposition of $\Gamma$ and $\Psi$, we can conclude that $\Psi=\Gamma^{-1}$.